Initially 100 milligrams of a radioactive . After 6 hours the mass had decayed by 3% the rate of decay is proportional to the amount of substance present at time t amount remaining after 24 hours .
1
Expert's answer
2021-07-09T14:09:31-0400

The rate of decay is proportional to the amount of substance present at time "t"


"\dfrac{dA}{dt}=kA"

"\dfrac{dA}{A}=kdt"

"\int\dfrac{dA}{A}=\int kdt"

"\ln |A|=kt+\ln C"

"A=Ce^{kt }"

Initially 100 milligrams of a radioactive


"100=Ce^{k(0)}=>C=100"

"A(t)=100e^{kt}"

After 6 hours the mass had decayed by 3%


"97=100e^{6k}"

"6k=\ln0.97"

"k=\dfrac{1}{6}\ln 0.97"

After 24 hours


"A(24)=100e^{{\ln 0.97 \over 6}(24)}"

"A(24)=100e^{4\ln 0.97}"

"A(24)=100(0.97)^4"

"A=88.529281\ mg"